Suppose a protein can exist in two conformations that have an energy difference of 2.0 kJ/mol. What is the estimated ratio of the conformations at 25 C? Which conformation is more abundant and by what factor? SHOW WORK!
Answer #1

Let the protein exist in two conformations 1 and 2 which exist in equilibrium as

Conformation 1 <======> Conformation 2

The equilibrium constant for the equilibrium is given as

K = [Conformation 2]/[Conformation 1]

The energy difference between the two conformations is 2.0 kJ/mol at 25ºC, i.e, ΔG0 = 2.0 kJ/mol at T = (25 + 273) K = 298 K.

We know that

ΔG0 = -RT*ln K

Plug in values and get

2.0 kJ/mol = -(8.314 J/mol.K)*(298 K)*ln K

====> (2.0 kJ/mol)*(1000 J)/(1 kJ) = -(8.314 J/mol.K)*(298 K)*ln K

====> 2000 J/mol = -(2477.572 J/mol)*ln K

====> ln K = (2000 J/mol)/(-2477.572 J/mol)

====> ln K = -0.80724

====> K = exp(-0.80724)

====> K = 0.44608 ≈ 0.45 (ans, correct to 2 sig. figs).

Therefore, the ratio of the two conformers at equilibrium is 0.45.

Since the ratio is less than 1, offcourse, conformation 1 is present in larger proportion as compared to conformation 2. It also turns out that conformation 1 is the more preferred conformation at equilibrium.